OpenStudy (anonymous):

triangle ABC has BC=1,and

OpenStudy (anonymous):

\[\sin^{23}\frac{ A }{ 2 }\cos^{48}\frac{ B }{ 2 }=\sin^{23} \frac{ B }{ 2 }\cos ^{48}\frac{ A }{ 2 }\] find AC

OpenStudy (anonymous):

startless-i dont know where to start

OpenStudy (anonymous):

\[AB^2=1+AC^2-2(AC)\cos A\] \[A+B+C=180\] \[\frac{ \sin A }{ AB }=\frac{\sin B }{1}=\frac{ \sin C }{ AC }\]

OpenStudy (anonymous):

@Callisto ,@dpaInc @ParthKohli

OpenStudy (anonymous):

@dpaInc

OpenStudy (anonymous):

Something's lacking, since the given trigonometry condition can be fulfilled by e.g. any isosceles triangle where ∠A = ∠B, hence AC is not determined. Sorry for late response.